# TAMO:Fine-Grained Root Cause Analysis via Tool-Assisted LLM Agent with Multi-Modality Observation Data

### v1

@@ Abstract
With the development of distributed systems, microservices and cloud native technologies have become central to modern enterprise software development. Despite bringing significant advantages, these technologies also increase system complexity and operational challenges. Traditional root cause analysis (RCA) struggles to achieve automated fault response, heavily relying on manual intervention. In recent years, large language models (LLMs) have made breakthroughs in contextual inference and domain knowledge integration, providing new solutions for Artificial Intelligence for Operations (AIOps). However, Existing LLM-based approaches face three key challenges: text input constraints, dynamic service dependency hallucinations, and context window limitations. To address these issues, we propose a tool-assisted LLM agent with multi-modality observation data, namely TAMO, for fine-grained RCA. It unifies multi-modal observational data into time-aligned representations to extract consistent features and employs specialized root cause localization and fault classification tools for perceiving the contextual environment. This approach overcomes the limitations of LLM in handling real-time changing service dependencies and raw observational data and guides LLM to generate repair strategies aligned with system contexts by structuring key information into a prompt. Experimental results show that TAMO performs well in root cause analysis when dealing with public datasets characterized by heterogeneity and common fault types, demonstrating its effectiveness.
@@ I Introduction
  • Context Limitations: Limited by fixed-length context windows, the LLM is unable to effectively process massive amounts of contextual monitoring data in real-time. This constraint often leads to data truncation or compression during analysis, making it difficult to utilize historical information to detect anomalous trends in system components, thereby affecting the rationale and accuracy of root cause analysis. So how to extract useful information from the complex system context and input it into the context-limited LLM is a key challenge.
  • Text Input Constraints: Existing large language models such as ChatGPT[1], DeepSeek[27], Qwen[3], etc., either only support text data as input or enable modality input by mapping different modal features to a unified textual semantic space. However, due to the lack of obvious correlation between time series and text, this transformation process often results in a significant loss of key feature information, severely hindering the analytical and reasoning capabilities of LLM. So how to efficiently unify multiple modalities of observation data to assist the LLM in providing accurate root cause analysis results is a key challenge.
  • Dynamic Dependence Hallucination: The LLM is based on static, pre-trained knowledge, which makes it challenging to model service dependencies in systems. It can only simulate simple graph structures using text prompts. As noted in [34], LLMs struggle with reasoning over complex graphs, which limits their ability to effectively capture service dependencies. Additionally, the dynamic nature of service call chains in the system means that evolving dependencies can cause the LLM to generate hallucinations that contradict actual dependencies, leading to inaccurate root cause analysis. So how to assist the LLM in modeling fault propagation paths using call chain relationships to achieve fine-grained root cause localization is a key challenge.
@@ III METHODOLOGY
To fully exploit the multimodal observability data of cloud native systems while integrating prior domain knowledge, we propose TAMO, Tool-Assisted LLM Agents that Integrates Multi-Modality Observation Data to address root cause entity localization and fault classification in large-scale system operations and maintenance. This workflow utilizes observation tools composed of a dual-branch collaborative diffusion model to effectively extract and integrate multimodal data from different entities. Additionally, task-specialized tools explicitly model the complex dependencies among entities using a causal graph, with feature propagation through the causal graph enabling accurate root cause analysis. Furthermore, to fully incorporate domain knowledge, we employ a Large Language Model (LLM) with operational expertise as an expert agent. The expert agent $\mathscr{A}$ integrates the outputs from the preceding tools $\mathscr{T}_{2}$-$\mathscr{T}_{3}$ with context knowledge of the system to offer human engineers a comprehensive fault analysis and recommended solutions. As shown in Fig.2, the proposed workflow consists of three tools and one agent, $\mathscr{T}_{1}$$\mathscr{T}_{3}$ and $\mathscr{A}$, each responsible for a specific task in the root cause analysis process:
@@
    RCA Expert Agent ($\mathscr{A}$): Leveraging the expertise of a Large Language Model (LLM), this agent synthesizes the outputs from previous tools along with system environment context to provide site reliability engineers (SREs) with a comprehensive fault analysis and recommended solutions.
@@ III-A Data Feature Observation Tool ($\mathscr{T}_{1}$)
$p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathcal{C}\right)$ $=\mathcal{N}\left(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t,\mathcal{C}\right),\mathbf{\Sigma}_{\theta}\left(\mathbf{x}_{t},t,\mathcal{C}\right)\right),$ (3)
$p_{\theta}\left(\mathbf{x}_{0:T},\mathcal{C}\right)$ $=p\left(\mathbf{x}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathcal{C}\right),$
@@
The original denoising diffusion model employs U-Net as the denoising backbone [17]. However, considering the characteristics of time-series data, we adopt a patch-based Transformer network [29] as the backbone for each branch to better capture the representations of time-series subsequences. The training objective of this model is to minimize the KL-divergence between the distributions $p$ and $q$. According to [17], the objective derivation is formulated as follows:
@@ III-B Root Cause Localization Tool ($\mathscr{T}_{2}$)
Through the dual-branch diffusion model of Tool $\mathscr{T}_{1}$, we obtain time-series data that integrates multimodal features. However, since this process operates at the level of system resource entities, the generated time-series $x^{\prime}_{ent}\in\mathbb{R}^{L\times d_{ent}}$ maintains the same dimensionality as the original resource entities, where $ent$ represents the corresponding resource entity, $L$ denotes the time window length, and $d_{ent}$ represents the feature dimension associated with each entity. Due to the heterogeneity of observability data across different resource entities in cloud-native systems, inconsistencies exist among them, which hinders unified learning in the model. To address this issue, we first perform data embedding on the features of different entities, mapping them into a unified dimensional space for consistency:
@@
$\mathcal{F}\left(H^{emb}\right)=\operatorname{FFT}\left(H^{emb}\right)=\sum_{i=0}^{L-1}H^{emb}_{i}e^{-j2\pi ki/K},\quad$ (10)
where $e^{-j2\pi ki/N}$ is the complex exponential Fourier basis function, $H^{emb}_{i}$ represents the time-series value corresponding to the time step $i$, and $k=0,1,\ldots,N-1$.
@@
$\operatorname{Indices}=\operatorname{argtopk}\left(\left|\mathcal{F}\left(H^{emb}\right)\right|,k=freq\_topk\right),$ (11)
@@
$\mathcal{F}_{\text{filtered }}=\mathcal{F}\left(H^{emb}\right)\odot Mask\in\mathbb{R}^{E\times K\times d_{model}},$ (13)
@@
$\hat{y}=\text{ MaxPool}\left(\mathbf{H}^{S}\right)\in\mathbb{R}^{E\times 1},$ (19)
@@ III-C Fault Classification Tool ($\mathscr{T}_{3}$)
$\text{Attention}=\text{softmax}\left(\frac{zW^{Q}\cdot z{W^{K}}^{T}}{\sqrt{d_{k}}}\right)zW^{V},$ (21)
where $W^{Q}$,$W^{K}$,$W^{V}$ are linear projection matrices. Subsequently, we concatenate the results of the multi-head attention to obtain the latent vector representation:
$V=\text{MultiHead}\left(zW^{Q},zW^{K},zW^{V}\right),$ (22)
@@
$\mathcal{L}=\mathcal{L}_{wBCE}+\beta\left||\mathbf{W}\right||^{2}_{2}.$ (25)
@@ III-D RCA Expert Agent ($\mathscr{A}$)
After the three tools $\mathscr{T}_{1}$-$\mathscr{T}_{3}$ have completed tasks such as data feature alignment, root cause localization, and fault classification, the workflow has achieved a relatively deep understanding of the fault. However, relying solely on these outputs makes it difficult to provide comprehensive and accurate failure analysis and remediation plans. Therefore, to further improve fault response and system recovery efficiency, we introduce the RCA expert agent $\mathscr{A}$. Using large language models (LLM), this agent synthesizes the predictions from the previous tools and incorporates the current context of cloud-native microservice architecture to offer a more comprehensive and accurate fault diagnosis and resolution strategy for site reliability engineers.
@@
Figure 3: Prompt templates for this RCA expert agent.
@@
AC@1 AC@3 AC@5 MiPr MaPr MiRe MaRe MiF1 MaF1
@@ IV-B2 Baselines
AC@1 AC@3 AC@5 MiPr MaPr MiRe MaRe MiF1 MaF1
@@ IV-B3 Evaluation metrics
TAMO is dedicated to locating root cause instances and identifying fault types. Based on previous research works[14],[23],[42],[15], we selected different evaluation metrics to evaluate the performance of model for these two tasks. Specifically, for the root cause localization task, we use top-k accuracy (AC@K) as the evaluation metric, which represents the probability that the true root cause entity appears within the top-k results produced by the method. Here, we set $k=\{1,3,5\}$. Given $S$ as the set of faults, $RC$ is the ground truth root cause, and $rc[k]$ are the top-k root causes generated by the model. Therefore, AC@K can be defined as follows:
$AC@k=\frac{1}{|S|}\sum_{i\in S}^{S}\begin{cases}1,&\text{if }RC_{i}\in rc_{i}[k]\\ 0,&\text{otherwise}\end{cases}$ (26)
@@ V Conclusion
This paper introduces an innovative Tool-Assisted LLM Agent framework named TAMO to address the complex challenges of root cause analysis in cloud-native microservice environments. By integrating domain-specific tools with large language models, TAMO effectively handles high-dimensional, multi-modal cloud-native observability data and accurately pinpoints fault origins amidst dynamically changing service dependencies. The approach not only overcomes the information loss issues when converting raw observability data into a format suitable for LLM processing but also addresses the challenge that static pre-trained knowledge struggles to adapt to real-time changes in service dependencies. Moreover, by structuring and distilling key results, it guides the LLM to generate contextually appropriate remediation strategies, thereby avoiding limitations imposed by fixed-length context windows. Our research demonstrates that using a dual-branch collaborative diffusion model enables capturing consistent patterns across different data modalities, providing a unified temporal representation for downstream analysis tasks. Additionally, the root cause localization method based on frequency-domain attention mechanisms enhances the accuracy of identifying causal dependency patterns while effectively filtering out low-frequency noise, achieving precise fault localization. The experimental results show that, compared to existing advanced methods and benchmarks, TAMO performs notably better on two standard datasets, highlighting its potential to improve fault diagnosis efficiency and accuracy. Further ablation studies and case analyses also validate the effectiveness of each component of the framework and the superior performance of their collaborative work. However, as system scale and complexity continue to grow, how to more efficiently process larger volumes of data streams and further enhance real-time interpretability and cross-modal semantic alignment remain critical areas for future research.

### v2

@@ Abstract
With the development of distributed systems, microservices and cloud native technologies have become central to modern enterprise software development. Despite bringing significant advantages, these technologies also increase system complexity and operational challenges. Implementing automated root cause analysis to reduce manual intervention has become a key aspect of modern software operations and maintenance. In recent years, large language models (LLMs) have made breakthroughs in root cause context inference and the integration of domain-specific knowledge, providing new solutions for Artificial Intelligence for IT Operations (AIOps). However, existing LLM-based approaches face three key challenges: context window limitation, modal input constraint, and dynamic dependence hallucination. To address these issues, we propose a tool-assisted LLM agent with multi-modality observation data, namely TAMO, for fine-grained RCA. It unifies multi-modal observational data into time-aligned representations to extract consistent features and employs specialized root cause localization and fault classification tools for perceiving the contextual environment. This approach overcomes the limitations of LLM in handling real-time changing service dependencies and raw observational data and guides LLM to generate repair strategies aligned with system contexts by structuring key information into a prompt. Experimental results show that TAMO performs well in root cause analysis when dealing with public datasets characterized by heterogeneity and common fault types, demonstrating its effectiveness.
@@ I Introduction
  • Context Window Limitation: Limited by fixed-length context windows, the LLM is unable to effectively process massive amounts of contextual monitoring data in real-time. This constraint often leads to data truncation or compression during analysis, making it difficult to utilize historical information to detect anomalous trends in system components, thereby affecting the rationale and accuracy of root cause analysis. So how to extract useful information from the complex system context and input it into the LLM with limited context length?
  • Modal Input Constraint: Existing large language models such as ChatGPT[1], DeepSeek[27], Qwen[3], etc., either only support text data as input or enable modality input by mapping different modal features to a unified textual semantic space. However, these models often suffer from significant key feature loss when dealing with observational data, especially metric data, due to the lack of a obvious correlation between time series data and semantic text, which hinders the ability to analyze and reason effectively. So how to efficiently unify multiple modalities of observation data to assist the LLM in providing accurate root cause analysis results?
  • Dynamic Dependence Hallucination: The LLM is based on static, pre-trained knowledge, which makes it challenging to model service dependencies in systems. It can only simulate simple graph structures using text prompts. As noted in [34], LLMs struggle with reasoning over complex graphs, which limits their ability to effectively capture service dependencies. Additionally, the dynamic nature of service call chains in the system means that evolving dependencies can cause the LLM to generate hallucination that contradict actual dependencies, leading to inaccurate root cause analysis. So how to assist the LLM in modeling fault propagation paths using call chain relationships to achieve fine-grained root cause localization?
@@ III METHODOLOGY
To fully exploit the multimodal observability data of cloud native systems while integrating prior domain knowledge, we propose TAMO, Tool-Assisted LLM Agent that Integrates Multi-Modality Observation Data to address root cause entity localization and fault classification in large-scale system operations and maintenance. This framework utilizes observation tools composed of a dual-branch collaborative diffusion model to effectively extract and integrate multimodal data from different entities. Additionally, task-specialized tools explicitly model the complex dependencies among entities using a causal graph, with feature propagation through the causal graph enabling accurate root cause analysis. Furthermore, to fully incorporate domain knowledge, we employ a LLM with operational expertise as an expert agent. The expert agent $\mathscr{A}$ integrates the outputs from the preceding tools $\mathscr{T}_{2}$-$\mathscr{T}_{3}$ with context knowledge of the system to offer human engineers a comprehensive fault analysis and recommended solutions. As shown in Fig.2, the proposed framework consists of three tools and one agent, $\mathscr{T}_{1}$$\mathscr{T}_{3}$ and $\mathscr{A}$, each responsible for a specific task in the root cause analysis process:
@@
    RCA Expert Agent ($\mathscr{A}$): Leveraging the expertise of a LLM, this agent synthesizes the outputs from previous tools along with system environment context to provide site reliability engineers (SREs) with a comprehensive fault analysis and recommended solutions.
@@ III-A Data Feature Observation Tool ($\mathscr{T}_{1}$)
$p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathcal{C}\right)$ $=\mathcal{N}\left(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t,\mathcal{C}\right),\mathbf{\Sigma}_{\theta}\left(\mathbf{x}_{t},t,\mathcal{C}\right)\right),$
$p_{\theta}\left(\mathbf{x}_{0:T},\mathcal{C}\right)$ $=p\left(\mathbf{x}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathcal{C}\right),$ (3)
@@
The original denoising diffusion model employs U-Net as the denoising backbone [17]. However, considering the characteristics of time-series data, we adopt a patch-based Transformer network [29] as the backbone for each branch to better capture the representations of time-series subsequences. The training objective of the denoising process is to minimize the KL-divergence between the distributions $p$ and $q$. According to [17], the objective derivation is formulated as follows:
@@
Through this denoising process, we obtain time series data $x^{\prime}_{ent}\in\mathbb{R}^{L\times d_{ent}}$ that integrates multimodal features for subsequent analysis, where $ent$ denotes the corresponding resource entity, $L$ denotes the length of the time window, and $d_{ent}$ denotes the dimension of the feature associated with each entity.
@@ III-B Root Cause Localization Tool ($\mathscr{T}_{2}$)
Through the dual-branch diffusion model of Tool $\mathscr{T}_{1}$, we obtain time-series data that integrates multimodal features. However, since this process operates at the level of system resource entities, the generated time-series $x^{\prime}_{ent}$ maintains the same dimensionality as the original resource entities. There exists heterogeneity in the observable data from different resource entities in cloud-native systems, leading to inconsistencies between them, which hinders the unified learning of the model. To address this issue, we first perform data embedding on the features of different entities, mapping them into a unified dimensional space for consistency:
@@
$\mathcal{F}\left(H^{emb}\right)=\operatorname{FFT}\left(H^{emb}\right)=\sum_{i=0}^{L-1}H^{emb}_{i}e^{-j2\pi ki/L},\quad$ (10)
where $e^{-j2\pi ki/N}$ is the complex exponential fourier basis function, $j$ is the imaginary unit, $H^{emb}_{i}$ represents the time-series value corresponding to the time step $i$, and $k=0,1,\ldots,N-1$ is the frequency index.
@@
$\operatorname{Indices}=topk\left(\left|\mathcal{F}\left(H^{emb}\right)\right|,k=freq\_topk\right),$ (11)
@@
$\mathcal{F}_{\text{filtered}}=\mathcal{F}\left(H^{emb}\right)\odot Mask,\mathcal{F}_{\text{filtered}}\in\mathbb{R}^{E\times K\times d_{model}},$ (13)
@@
$\hat{y}=\text{ MaxPool}\left(\mathbf{H}^{S}\right)\in\mathbb{R}^{E},$ (19)
@@ III-C Fault Classification Tool ($\mathscr{T}_{3}$)
$\text{Attention}=\text{softmax}\left(\frac{zW^{\hat{Q}}\cdot z{W^{\hat{K}}}^{T}}{\sqrt{d_{k}}}\right)zW^{\hat{V}},$ (21)
where $W^{\hat{Q}}$,$W^{\hat{K}}$,$W^{\hat{V}}$ are linear projection matrices. Subsequently, we concatenate the results of the multi-head attention to obtain the latent vector representation:
$V=\text{MultiHead}\left(zW^{\hat{Q}},zW^{\hat{K}},zW^{\hat{V}}\right),$ (22)
@@
$\mathcal{L}=\mathcal{L}_{wBCE}+\beta\|\mathbf{W}\|^{2}_{2}.$ (25)
@@
Figure 3: Prompt templates for this RCA expert agent.
@@ III-D RCA Expert Agent ($\mathscr{A}$)
After the three tools $\mathscr{T}_{1}$-$\mathscr{T}_{3}$ have completed tasks such as data feature alignment, root cause localization, and fault classification, the framework has achieved a relatively deep understanding of the fault. However, relying solely on these outputs makes it difficult to provide comprehensive and accurate failure analysis and remediation plans. Therefore, to further improve fault response and system recovery efficiency, we introduce the RCA expert agent $\mathscr{A}$. Using large language models, this agent synthesizes the predictions from the previous tools and incorporates the current context of cloud-native microservice architecture to offer a more comprehensive and accurate fault diagnosis and resolution strategy for site reliability engineers.
@@
Acc@1 Acc@3 Acc@5 MiPr MaPr MiRe MaRe MiF1 MaF1
@@ IV-B2 Baselines
Acc@1 Acc@3 Acc@5 MiPr MaPr MiRe MaRe MiF1 MaF1
@@ IV-B3 Evaluation metrics
TAMO is dedicated to locating root cause instances and identifying fault types. Based on previous research works[14],[23],[42],[15], we selected different evaluation metrics to evaluate the performance of model for these two tasks. Specifically, for the root cause localization task, we use top-k accuracy (Acc@K) as the evaluation metric, which represents the probability that the true root cause entity appears within the top-k results produced by the method. Here, we set $k=\{1,3,5\}$. Given $S$ as the set of faults, $RC$ is the ground truth root cause, and $rc[k]$ are the top-k root causes generated by the model. Therefore, Acc@K can be defined as follows:
$Acc@k=\frac{1}{|S|}\sum_{i\in S}^{S}\begin{cases}1,&\text{if }RC_{i}\in rc_{i}[k]\\ 0,&\text{otherwise}\end{cases}$ (26)
@@ V Conclusion
This paper introduces an innovative tool-assisted LLM agent framework named TAMO to address the complex challenges of root cause analysis in cloud-native microservice environments. By integrating domain-specific tools with large language models, TAMO efficiently handles high-dimensional, multi-modal cloud-native observability data and accurately locates root causes of failures in dynamically changing service dependencies. The approach not only overcomes the information loss issues when converting raw observability data into a format suitable for LLM processing but also addresses the challenge that static pre-trained knowledge struggles to adapt to real-time changes in service dependencies. The experimental results show that, compared to existing advanced methods and benchmarks, TAMO performs notably better on two standard datasets, highlighting its potential to improve fault diagnosis efficiency and accuracy. However, as system scale and complexity continue to grow, how to more efficiently process larger volumes of data streams and further enhance real-time interpretability and cross-modal semantic alignment remain critical areas for future research.